# F5-TTS: A Fairytaler that Fakes Fluent and Faithful Speech with Flow Matching

### v2

@@ Abstract
This paper introduces F5-TTS, a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer (DiT). Without requiring complex designs such as duration model, text encoder, and phoneme alignment, the text input is simply padded with filler tokens to the same length as input speech, and then the denoising is performed for speech generation, which was originally proved feasible by E2 TTS. However, the original design of E2 TTS makes it hard to follow due to its slow convergence and low robustness. To address these issues, we first model the input with ConvNeXt to refine the text representation, making it easy to align with the speech. We further propose an inference-time Sway Sampling strategy, which significantly improves our model’s performance and efficiency. This sampling strategy for flow step can be easily applied to existing flow matching based models without retraining. Our design allows faster training and achieves an inference RTF of 0.15, which is greatly improved compared to state-of-the-art diffusion-based TTS models. Trained on a public 100K hours multilingual dataset, our Fairytaler Fakes Fluent and Faithful speech with Flow matching (F5-TTS) exhibits highly natural and expressive zero-shot ability, seamless code-switching capability, and speed control efficiency. Demo samples can be found at https://SWivid.github.io/F5-TTS. We release all code and checkpoints to promote community development [^1] https://github.com/SWivid/F5-TTS.
@@ 1 Introduction
Recent research in Text-to-Speech (TTS) has experienced great advancement [1, 2, 3, 4, 5, 6, 7, 8]. With a few seconds of audio prompt, current TTS models are able to synthesize speech for any given text and mimic the speaker of audio prompt [9, 10]. The synthesized speech can achieve high fidelity and naturalness that they are almost indistinguishable from human speech [11, 12, 13, 14].
While autoregressive (AR) based TTS models exhibit an intuitive way of consecutively predicting the next token(s) and have achieved promising zero-shot TTS capability, the inherent limitations of AR modeling require extra efforts addressing issues such as inference latency and exposure bias [15, 16, 17, 18, 19]. Moreover, the quality of speech tokenizer is essential for AR models to achieve high-fidelity synthesis [20, 21, 22, 23, 24, 25, 26]. Thus, there have been studies exploring direct modeling in continuous space [27, 28, 29] to enhance synthesized speech quality recently.
Although AR models demonstrate impressive zero-shot performance as they perform implicit duration modeling and can leverage diverse sampling strategies, non-autoregressive (NAR) models benefit from fast inference through parallel processing, and effectively balance synthesis quality and latency. Notably, diffusion models [30, 31] contribute most to the success of current NAR speech models [11, 12]. In particular, Flow Matching with Optimal Transport path (FM-OT) [32] is widely used in recent research fields not only text-to-speech [14, 33, 34, 35, 36] but also image generation [37] and music generation [38].
@@
Figure 1: An overview of F5-TTS training (left) and inference (right). The model is trained on the text-guided speech-infilling task and condition flow matching loss. The input text is converted to a character sequence, padded with filler tokens to the same length as input speech, and refined by ConvNeXt blocks before concatenation with speech input. The inference leverages Sway Sampling for flow steps, with the model and an ODE solver to generate speech from sampled noise.
@@
Unlike AR-based models, the alignment modeling between input text and synthesized speech is crucial and challenging for NAR-based models. While NaturalSpeech 3 [12] and Voicebox [14] use frame-wise phoneme alignment; Matcha-TTS [34] adopts monotonic alignment search and relies on phoneme-level duration model; recent works find that introducing such rigid and inflexible alignment between text and speech hinders the model from generating results with higher naturalness [36, 39].
@@
E3 TTS [40] abandons phoneme-level duration and applies cross-attention on the input sequence but yields limited audio quality. DiTTo-TTS [35] uses Diffusion Transformer (DiT) [41] with cross-attention conditioned on encoded text from a pretrained language model. To further enhance alignment, it uses the pretrained language model to finetune the neural audio codec, infusing semantic information into the generated representations. In contrast, E2 TTS [36], based on Voicebox [14], adopts a simpler way, which removes the phoneme and duration predictor and directly uses characters padded with filler tokens to the length of mel spectrograms as input. This simple scheme also achieves very natural and realistic synthesized results. However, we found that robustness issues exist in E2 TTS for the text and speech alignment. Seed-TTS [39] employs a similar strategy and achieves excellent results, though not elaborated in model details. In these ways of not explicitly modeling phoneme-level duration, models learn to assign the length of each word or phoneme according to the given total sequence length, resulting in improved prosody and rhythm.
@@
In this paper, we propose F5-TTS, a Fairytaler that Fakes Fluent and Faithful speech with Flow matching. Maintaining the simplicity of pipeline without phoneme alignment, duration predictor, text encoder, and semantically infused codec model, F5-TTS leverages the Diffusion Transformer with ConvNeXt V2 [42] to better tackle text-speech alignment during in-context learning. We stress the deep entanglement of semantic and acoustic features in the E2 TTS model design, which has inherent problems and will pose alignment failure issues that could not simply be solved with re-ranking. With in-depth ablation studies, our proposed F5-TTS demonstrates stronger robustness, in generating more faithful speech to the text prompt, while maintaining comparable speaker similarity. Additionally, we introduce an inference-time sampling strategy for flow steps substantially improving naturalness, intelligibility, and speaker similarity of generation. This approach can be seamlessly integrated into existing flow matching based models without retraining.
@@ 2.1 Flow Matching
$\mathcal{L}_{FM}(\theta)=E_{t,p_{t}(x)}\left\|v_{t}(x)-u_{t}(x)\right\|^{2},$ (1)
where $\theta$ parameterizes the neural network, $t\sim\mathcal{U}[0,1]$ and $x\sim p_{t}(x)$. The model $v_{t}$ is trained over the entire flow step and data range, ensuring it learns to handle the entire transformation process from the initial distribution to the target distribution.
As we have no prior knowledge of how to approximate $p_{t}$ and $u_{t}$, a conditional probability path $p_{t}(x|x_{1})=\mathcal{N}(x\ |\ \mu_{t}(x_{1}),\sigma_{t}(x_{1})^{2}I)$ is considered in actual training, and the Conditional Flow Matching (CFM) loss is proved to have identical gradients w.r.t. $\theta$ [32]. $x_{1}$ is the random variable corresponding to training data. $\mu$ and $\sigma$ is the time-dependent mean and scalar standard deviation of Gaussian distribution.
@@
$\mathcal{L}_{\text{CFM}}(\theta)=E_{t,q(x_{1}),p(x_{0})}\|v_{t}(\psi_{t}(x_{0}))-\frac{d}{dt}\psi_{t}(x_{0})\|^{2}.$ (2)
Further leveraging Optimal Transport form $\psi_{t}(x)=(1-t)x+tx_{1}$, we have the OT-CFM loss,
$\mathcal{L}_{\text{CFM}}(\theta)=E_{t,q(x_{1}),p(x_{0})}\|v_{t}((1-t)x_{0}+tx_{1})-(x_{1}-x_{0})\|^{2}.$ (3)
To view in a more general way [43], if formulating the loss in terms of log signal-to-noise ratio (log-SNR) $\lambda$ instead of flow step $t$, and parameterizing to predict $x_{0}$ ($\epsilon$, commonly stated in diffusion model) instead of predict $x_{1}-x_{0}$, the CFM loss is equivalent to the $v$-prediction [44] loss with cosine schedule.
For inference, given sampled noise $x_{0}$ from initial distribution $p_{0}$, flow step $t\in[0,1]$ and condition with respect to generation task, the ordinary differential equation (ODE) solver [45] is used to evaluate $\psi_{1}(x_{0})$ the integration of $d\psi_{t}(x_{0})/dt$ with $\psi_{0}(x_{0})=x_{0}$. The number of function evaluations (NFE) is the times going through the neural network as we may provide multiple flow step values from 0 to 1 as input to approximate the integration. Higher NFE will produce more accurate results and certainly take more calculation time.
@@ 2.2 Classifier-Free Guidance
Classifier Guidance (CG) is proposed by [46], functions by adding the gradient of an additional classifier, while such an explicit way to condition the generation process may have several problems. Extra training of the classifier is required and the generation result is directly affected by the quality of the classifier. Adversarial attacks might also occur as the guidance is introduced through the way of updating the gradient. Thus deceptive images with imperceptible details to human eyes may be generated, which are not conditional.
Classifier-Free Guidance (CFG) [47] proposes to replace the explicit classifier with an implicit classifier without directly computing the explicit classifier and its gradient. The gradient of a classifier can be expressed as a combination of conditional generation probability and unconditional generation probability. By dropping the condition with a certain rate during training, and linear extrapolating the inference outputs with and without condition $c$, the final guided result is obtained. We could balance between fidelity and diversity of the generated samples with
$v_{t,CFG}=v_{t}(\psi_{t}(x_{0}),c)+\alpha(v_{t}(\psi_{t}(x_{0}),c)-v_{t}(\psi_{t}(x_{0})))$ (4)
@@ 3 Method
This work aims to build a high-level text-to-speech synthesis system. We trained our model on the text-guided speech-infilling task [48, 14]. Based on recent research [35, 36, 49], it is promising to train without phoneme-level duration predictor and can achieve higher naturalness in zero-shot generation deprecating explicit phoneme-level alignment. We adopt a similar pipeline as E2 TTS [36] and propose our advanced architecture F5-TTS, addressing the slow convergence (timbre learned well at an early stage but struggled to learn alignment) and robustness issues (failures on hard case generation) of E2 TTS. We also propose a Sway Sampling strategy for flow steps at inference, which significantly improves our model’s performance in faithfulness to reference text and speaker similarity.
@@ 3.1 Pipeline
Following E2 TTS, we directly use alphabets and symbols for English. We opt for full pinyin to facilitate Chinese zero-shot generation. By breaking the raw text into such character sequence and padding it with filler tokens to the same length as mel frames, we form an extended sequence $z$ with $c_{i}$ denoting the $i$-th character:
@@
See from Fig.1 (right), we start from a sampled noise $x_{0}$, and what we want is the other end of flow $x_{1}$. Thus we use the ODE solver to gradually integrate from $\psi_{0}(x_{0})=x_{0}$ to $\psi_{1}(x_{0})=x_{1}$, given $d\psi_{t}(x_{0})/dt=v_{t}(\psi_{t}(x_{0}),x_{ref},z_{ref\cdot gen})$. During inference, the flow steps are provided in an ordered way, $e.g.$, uniformly sampled a certain number from 0 to 1 according to the NFE setting.
After getting the generated mel with model $v_{t}$ and ODE solver, we discard the part of $x_{ref}$. Then we leverage a vocoder to convert the mel back to speech signal.
@@ 3.2 F5-TTS
E2 TTS directly concatenates the padded character sequence with input speech, thus deeply entangling semantic and acoustic features with a large length gap of effective information, which is the underlying cause of hard training and poses several problems in a zero-shot scenario (Sec.5.1). To alleviate the problem of slow convergence and low robustness, we propose F5-TTS which accelerates training and inference and shows a strong robustness in generation. Also, an inference-time Sway Sampling is introduced, which allows inference faster (using less NFE) while maintaining performance. This sampling way of flow step can be directly applied to other CFM models.
Model As shown in Fig.1, we use latent Diffusion Transformer (DiT) [41] as backbone. To be specific, we use DiT blocks with zero-initialized adaptive Layer Norm (adaLN-zero). To enhance the model’s alignment ability, we also leverage ConvNeXt V2 blocks [42]. Its predecessor ConvNeXt V1 [50] is used in many works and shows a strong temporal modeling capability in speech domain tasks [51, 52].
As described in Sec.3.1, the model input is character sequence, noisy speech, and masked speech. Before concatenation in the feature dimension, the character sequence first goes through ConvNeXt blocks. Experiments have shown that this way of providing individual modeling space allows text input to better prepare itself before later in-context learning. Unlike the phoneme-level force alignment done in Voicebox, a rigid boundary for text is not explicitly introduced. The semantic and acoustic features are jointly learned with the entire model. And unlike the way of feeding the model with inputs of significant length difference (length with effective information) as E2 TTS does, our design mitigates such gap.
@@
We adopt some position embedding settings in Voicebox. The flow step is embedded with a sinusoidal position. The concatenated input sequence is added with a convolutional position embedding. We apply a rotary position embedding (RoPE) [53] for self-attention rather than symmetric bi-directional ALiBi bias [54]. And for extended character sequence $\hat{y}$, we also add it with an absolute sinusoidal position embedding before feeding it into ConvNeXt blocks.
Compared with Voicebox and E2 TTS, we abandoned the U-Net [55] style skip connection structure and switched to using DiT with adaLN-zero. Without a phoneme-level duration predictor and explicit alignment process, and nor with extra text encoder and semantically infused neural codec model in DiTTo-TTS, we give the text input a little freedom (individual modeling space) to let it prepare itself before concatenation and in-context learning with speech input.
Sampling As stated in Sec.2.1, the CFM could be viewed as v-prediction with a cosine schedule. For image synthesis, [37] propose to further schedule the flow step with a single-peak logit-normal [56] sampling, in order to give more weight to intermediate flow steps by sampling them more frequently. We speculate that such sampling distributes the model’s learning difficulty more evenly over different flow step $t\in[0,1]$.

### v3

@@ Abstract
This paper introduces F5-TTS, a fully non-autoregressive text-to-speech system based on flow matching with Diffusion Transformer (DiT). Without requiring complex designs such as duration model, text encoder, and phoneme alignment, the text input is simply padded with filler tokens to the same length as input speech, and then the denoising is performed for speech generation, which was originally proved feasible by E2 TTS. However, the original design of E2 TTS makes it hard to follow due to its slow convergence and low robustness. To address these issues, we first model the input with ConvNeXt to refine the text representation, making it easy to align with the speech. We further propose an inference-time Sway Sampling strategy, which significantly improves our model’s performance and efficiency. This sampling strategy for flow step can be easily applied to existing flow matching based models without retraining. Our design allows faster training and achieves an inference RTF of 0.15, which is greatly improved compared to state-of-the-art diffusion-based TTS models. Trained on a public 100K hours multilingual dataset, our F5-TTS exhibits highly natural and expressive zero-shot ability, seamless code-switching capability, and speed control efficiency. We have released all codes and checkpoints to promote community development, at https://SWivid.github.io/F5-TTS/.
F5-TTS: A Fairytaler that Fakes Fluent and Faithful Speech
with Flow Matching
Yushen Chen1, Zhikang Niu1, Ziyang Ma1, Keqi Deng2 Chunhui Wang3, Jian Zhao3, Kai Yu1, Xie Chen1thanks: Corresponding author 1MoE Key Lab of Artificial Intelligence, X-LANCE Lab, Shanghai Jiao Tong University 2University of Cambridge 3Geely Automobile Research Institute (Ningbo) Company Ltd. {swivid,chenxie95}@sjtu.edu.cn
@@ 1 Introduction
Figure 1: An overview of F5-TTS training (left) and inference (right). The model is trained on the text-guided speech-infilling task and condition flow matching loss. The input text is converted to a character sequence, padded with filler tokens to the same length as input speech, and refined by ConvNeXt V2 blocks before concatenation with speech input. The inference leverages Sway Sampling for flow steps, with the model and an ODE solver to generate speech from sampled noise.
Recent research in Text-to-Speech (TTS) has experienced great advancement (Shen et al., 2018; Li et al., 2019; Ren et al., 2020; Kim et al., 2020, 2021; Popov et al., 2021; Wang et al., 2023b; Tan et al., 2024). With a few seconds of audio prompt, current TTS models are able to synthesize speech for any given text and mimic the speaker of audio prompt (Wang et al., 2023a; Zhang et al., 2023b). The synthesized speech can achieve high fidelity and naturalness that they are almost indistinguishable from human speech (Shen et al., 2023; Ju et al., 2024; Chen et al., 2024; Le et al., 2024).
While autoregressive (AR) based TTS models exhibit an intuitive way of consecutively predicting the next token(s) and have achieved promising zero-shot TTS capability, the inherent limitations of AR modeling require extra efforts addressing issues such as inference latency and exposure bias (Song et al., 2024; Du et al., 2024a; Han et al., 2024; Xin et al., 2024; Peng et al., 2024). Moreover, the quality of speech tokenizer is essential for AR models to achieve high-fidelity synthesis (Zeghidour et al., 2021; Défossez et al., 2022; Wu et al., 2023; Yang et al., 2023; Zhang et al., 2023a; Bai et al., 2024; Niu et al., 2024). Thus, there have been studies exploring direct modeling in continuous space (Liu et al., 2024a; Li et al., 2024; Meng et al., 2024) to enhance synthesized speech quality recently.
Although AR models demonstrate impressive zero-shot performance as they perform implicit duration modeling and can leverage diverse sampling strategies, non-autoregressive (NAR) models benefit from fast inference through parallel processing, and effectively balance synthesis quality and latency. Notably, diffusion models (Ho et al., 2020; Song et al., 2020) contribute most to the success of current NAR speech models (Shen et al., 2023; Ju et al., 2024). In particular, Flow Matching with Optimal Transport path (FM-OT) (Lipman et al., 2022) is widely used in recent research fields not only text-to-speech (Le et al., 2024; Guo et al., 2024b; Mehta et al., 2024; Lee et al., 2024; Eskimez et al., 2024) but also image generation (Esser et al., 2024) and music generation (Fei et al., 2024).
@@
Unlike AR-based models, the alignment modeling between input text and synthesized speech is crucial and challenging for NAR-based models. While NaturalSpeech 3 (Ju et al., 2024) and Voicebox (Le et al., 2024) use frame-wise phoneme alignment; Matcha-TTS (Mehta et al., 2024) adopts monotonic alignment search (Kim et al., 2020) and relies on a phoneme-level duration model; recent works find that introducing such rigid alignment between text and speech hinders the model from generating results with higher naturalness (Eskimez et al., 2024; Anastassiou et al., 2024).
@@
E3 TTS (Gao et al., 2023a) abandons phoneme-level duration and applies cross-attention on the input sequence but yields limited audio quality. DiTTo-TTS (Lee et al., 2024) uses Diffusion Transformer (DiT) (Peebles and Xie, 2023) with cross-attention conditioned on encoded text from a pretrained language model. To further enhance alignment, it uses the pretrained language model to finetune the neural audio codec, infusing semantic information into the generated representations. In contrast, E2 TTS (Eskimez et al., 2024), based on Voicebox (Le et al., 2024), adopts a simpler way, which removes the phoneme and duration predictor and directly uses characters padded with filler tokens to the length of mel spectrograms as input. This simple scheme also achieves very natural and realistic synthesized results. However, we found that robustness issues exist in E2 TTS for the text and speech alignment. Seed-TTS (Anastassiou et al., 2024) employs a similar strategy and achieves excellent results, though not elaborated in model details. In these ways of not explicitly modeling phoneme-level duration, models learn to assign the length of each word or phoneme according to the given total sequence length, resulting in improved prosody and rhythm.
@@
In this paper, we propose F5-TTS, a Fairytaler that Fakes Fluent and Faithful speech with Flow matching. Maintaining the simplicity of pipeline without phoneme alignment, duration predictor, text encoder, and semantically infused codec model, F5-TTS leverages the Diffusion Transformer with ConvNeXt V2 (Woo et al., 2023) to better tackle text-speech alignment during in-context learning. We stress the deep entanglement of semantic and acoustic features in the E2 TTS model design, which has inherent problems and will pose alignment failure issues that could not simply be solved with re-ranking. With in-depth ablation studies, our proposed F5-TTS demonstrates stronger robustness, in generating more faithful speech to the text prompt, while maintaining comparable speaker similarity. Additionally, we introduce an inference-time sampling strategy for flow steps substantially improving naturalness, intelligibility, and speaker similarity of generation. This approach can be seamlessly integrated into existing flow matching based models without retraining.
@@ 2.1 Flow Matching
$\mathcal{L}_{\text{FM}}(\theta)=E_{t,p_{t}(x)}\left\|v_{t}(x)-u_{t}(x)\right\|^{2},$
where $\theta$ parameterizes the neural network, $t\sim\mathcal{U}[0,1]$ and $x\sim p_{t}(x)$. $v_{t}$ is trained over the entire flow step and data range, ensuring it learns to handle the entire transformation process from the initial distribution to the target distribution.
As we have no prior knowledge of how to approximate $p_{t}$ and $u_{t}$, a conditional probability path $p_{t}(x|x_{1})=\mathcal{N}(x\ |\ \mu_{t}(x_{1}),\sigma_{t}(x_{1})^{2}I)$ is considered in actual training, and the Conditional Flow Matching (CFM) loss is proved to have identical gradients w.r.t. $\theta$ (Lipman et al., 2022). $x_{1}$ is the random variable corresponding to training data. $\mu$ and $\sigma$ is the time-dependent mean and scalar standard deviation of Gaussian distribution.
@@
$\mathcal{L}_{\text{CFM}}(\theta)=E_{t,q(x_{1}),p(x_{0})}\|v_{t}(\psi_{t}(x_{0}))-\frac{d}{dt}\psi_{t}(x_{0})\|^{2}.$
Further leveraging Optimal Transport (OT) form $\psi_{t}(x)=(1-t)x+tx_{1}$, the OT-CFM loss is
$\mathcal{L}_{\text{CFM}}(\theta)=E_{t,q(x_{1}),p(x_{0})}\|v_{t}((1-t)x_{0}+tx_{1})-(x_{1}-x_{0})\|^{2}.$
To view in a more general way (Kingma and Gao, 2024), if formulating the loss in terms of log signal-to-noise ratio (log-SNR) $\lambda$ instead of flow step $t$, and parameterizing to predict $x_{0}$ ($\epsilon$, commonly stated in diffusion model) instead of predict $x_{1}-x_{0}$, the CFM loss is equivalent to the v-prediction (Salimans and Ho, 2022) loss with cosine schedule.
For inference, given sampled noise $x_{0}$ from initial distribution $p_{0}$, flow step $t\in[0,1]$ and condition with respect to generation task, the ordinary differential equation (ODE) solver (Chen, 2018) is used to evaluate $\psi_{1}(x_{0})$ the integration of $d\psi_{t}(x_{0})/dt$ with $\psi_{0}(x_{0})=x_{0}$. The number of function evaluations (NFE) is the times going through the neural network as we may provide multiple flow step values from 0 to 1 as input to approximate the integration. Higher NFE will produce more accurate results and certainly take more calculation time.
@@ 2.2 Classifier-Free Guidance
Classifier Guidance (CG) is proposed by Dhariwal and Nichol (2021), functions by adding the gradient of an additional classifier, while such an explicit way to condition the generation process may have several problems. Extra training of the classifier is required and the generation result is directly affected by the quality of the classifier. Adversarial attacks might also occur as the guidance is introduced through the way of updating the gradient. Thus deceptive images with imperceptible details to human eyes may be generated, which are not conditional.
Classifier-Free Guidance (CFG) (Ho and Salimans, 2022) proposes to replace the explicit classifier with an implicit classifier without directly computing the explicit classifier and its gradient. The gradient of a classifier can be expressed as a combination of conditional generation probability and unconditional generation probability. By dropping the condition with a certain rate during training, and linear extrapolating the inference outputs with and without condition $c$, the final guided result is obtained. We could balance between fidelity and diversity of the generated samples with
$v_{t,\text{CFG}}=v_{t}(\psi_{t}(x_{0}),c)+\alpha(v_{t}(\psi_{t}(x_{0}),c)-v_{t}(\psi_{t}(x_{0})))$
@@ 3 Method
This work aims to build a high-level text-to-speech synthesis system. We trained our model on the text-guided speech-infilling task (Bai et al., 2022; Le et al., 2024). Based on recent research (Lee et al., 2024; Eskimez et al., 2024; Liu et al., 2024b), it is promising to train without phoneme-level duration predictor and can achieve higher naturalness in zero-shot generation deprecating explicit phoneme-level alignment. We propose our advanced architecture with faster convergence and more robust generation. We also propose an inference-time flow step sampling strategy, which significantly improves our model’s performance in faithfulness to reference text and speaker similarity.
@@ 3.1 Pipeline
Following E2 TTS, we directly use alphabets and symbols for English. We opt for full pinyin to facilitate Chinese zero-shot generation. By breaking the raw text into such character sequence and padding it with filler tokens $\langle F\rangle$ to the same length as mel frames, we form an extended sequence $z$ with $c_{i}$ denoting the $i$-th character:
@@
See from Fig.1 (right), we start from a sampled noise $x_{0}$, and what we want is the other end of flow $x_{1}$. Thus we use the ODE solver to gradually integrate from $\psi_{0}(x_{0})=x_{0}$ to $\psi_{1}(x_{0})=x_{1}$, given $d\psi_{t}(x_{0})/dt=v_{t}(\psi_{t}(x_{0}),x_{ref},z_{ref\cdot gen})$. During inference, the flow steps are provided in an ordered way, $e.g.$, uniformly sampled a certain number from 0 to 1 according to the NFE setting. After getting the generated mel with model $v_{t}$ and ODE solver, we discard the part of $x_{ref}$. Then we leverage a vocoder to convert the mel back to waveform.
@@ 3.2 F5-TTS
E2 TTS directly concatenates the padded character sequence with input speech sequence, deeply entangling semantic and acoustic features with a large length gap of effective information, which is the underlying cause of hard training and poses several problems in a zero-shot scenario (Sec.5.1). To alleviate the problem of slow convergence and low robustness, we propose F5-TTS which accelerates training and inference and shows a strong robustness in generation. Also, an inference-time Sway Sampling is introduced, which allows inference faster (using less NFE) while maintaining performance. This sampling way of flow step can be directly applied to other CFM-based models without retraining.
Model As shown in Fig.1, we use latent Diffusion Transformer (DiT) (Peebles and Xie, 2023) as backbone. To be specific, we use DiT blocks with zero-initialized adaptive Layer Norm (adaLN-zero). To enhance the model’s alignment ability, we also leverage ConvNeXt V2 blocks (Woo et al., 2023). Its predecessor ConvNeXt V1 (Liu et al., 2022) is used in many works and shows a strong temporal modeling capability in speech domain tasks (Siuzdak, 2023; Okamoto et al., 2024).
As described in Sec.3.1, the model input is character sequence, noisy speech, and masked speech. Before concatenation in the feature dimension, the character sequence first goes through ConvNeXt blocks. Experiments have shown that this way of providing individual modeling space allows text input to better prepare itself before later in-context learning. Unlike the phoneme-level force alignment done in Voicebox, a rigid boundary for text is not explicitly introduced. The semantic and acoustic features are jointly learned with the entire model. Not directly feeding the model with inputs of significant length gap as E2 TTS does, the proposed text refinement mitigates the impact of using inputs with mismatched effective information lengths, despite equal physical length in magnitude as E2 TTS.
@@
We adopt some position embedding settings in Voicebox. The flow step is embedded with a sinusoidal position. The concatenated input sequence is added with a convolutional position embedding. We apply a rotary position embedding (RoPE) (Su et al., 2024) for self-attention rather than symmetric bi-directional ALiBi bias (Press et al., 2021). And for extended character sequence $z$, we also add it with an absolute sinusoidal position embedding before feeding it into ConvNeXt blocks.
Compared with Voicebox and E2 TTS, we abandoned the U-Net (Ronneberger et al., 2015) style skip connection structure and switched to using DiT with adaLN-zero. Without a phoneme-level duration predictor and explicit alignment process, and nor with extra text encoder and semantically infused neural codec model in DiTTo-TTS, we give the text input a little freedom (individual modeling space) to let it prepare itself before concatenation and in-context learning with speech input.
Sampling As stated in Sec.2.1, the CFM could be viewed as v-prediction with a cosine schedule. For image synthesis, Esser et al. (2024) propose to further schedule the flow step with a single-peak logit-normal (Atchison and Shen, 1980) sampling, in order to give more weight to intermediate flow steps by sampling them more frequently. We speculate that such sampling distributes the model’s learning difficulty more evenly over different flow step $t\in[0,1]$.